# Discarding Lavrentiev’s Gap in Non-automonous and Non-Convex Variational Problems

### v1

@@ 4 Between balance conditions ($H^{\textrm{iso}}$) and ($H^{\textrm{conv}}$)
In particular, ${w}(s)+s^{\max(p,N)}\leq{\mathcal{C}}_{\widehat{\mathcal{k}}}\mathcal{k}_{2}{\varepsilon}^{-N}+{\mathcal{C}}_{\widehat{\mathcal{k}}}$. Without loss of generality, we may assume that ${\varepsilon}\leq\text{diam}(\Omega)$ and that $\mathcal{k}_{2}$ is sufficiently large to ensure that $\mathcal{k}_{2}\text{diam}(\Omega)^{-N}\geq 1$. Therefore, we have ${w}(s)+s^{\max(p,N)}\leq 2{\mathcal{C}}_{\widehat{\mathcal{k}}}\mathcal{k}_{2}{\varepsilon}^{-N}$. Denoting $\widetilde{\mathcal{k}}=(\mathcal{k}_{1},2{\mathcal{C}}_{\widehat{\mathcal{k}}}\mathcal{k}_{2})$, we get from ($H^{\textrm{iso}}$) that ${w}_{x}(s)\leq{\mathcal{C}}_{\widetilde{\mathcal{k}}}\left({w}(s)+1\right)$, which by (4.4) yields
@@
which is ($H^{\textrm{conv}}$) for $f$ with $\widetilde{\mathcal{C}}_{\mathcal{k}}=\left({\mathcal{C}}_{\widetilde{\mathcal{k}}}{\mathcal{C}}_{\widehat{\mathcal{k}}}+{\mathcal{C}}_{\widetilde{\mathcal{k}}}+{\mathcal{C}}_{\widehat{\mathcal{k}}}\right)$.
@@ 5.2 The convex extension $\widehat{E}$ of the energy $E$
In order to define $\widehat{E}_{\Omega\times{\mathbb{R}}}$, we need to introduce the $\widehat{\ }$-operation, which goes back to [Dal-Maso] in the context of integral representation of $\Gamma$-limits of variational integrals. For any non-negative convex function $h:{\mathbb{R}}^{N}\to{[0,\infty)}$, we define its recession function $h^{\infty}:{\mathbb{R}^{N}}\to[0,\infty]$ as
@@
It follows from the convexity of $h$ that $\widehat{h}$ is convex on ${\mathbb{R}}^{N+1}$ and positively homogeneous of degree $1$, see e.g. [Bousquet-Pisa, Lemma 8.1]. In particular, for $h\equiv{1}$, we get for every $q=(q^{x},q^{t})\in{\mathbb{R}}^{N}\times{\mathbb{R}}$ with $q^{t}\leq 0$,
@@
Applying the operation $\ \widehat{}$ to $h=f(x,t,\cdot)$ for a.e. $x\in\Omega$ and every $t\in{\mathbb{R}}$, we obtain a map $q\mapsto\widehat{f(x,t,\cdot)}(q)$. In order to simplify the notation, we write
@@
$\widehat{f}(x,t,q^{x},q^{t})=\begin{cases}-q^{t}f(x,t,\frac{q^{x}}{-q^{t}})\,,&\textrm{ if }q^{t}<0\,,\\ \infty\,,\quad&\textrm{ if $q^{t}>0$ or $q^{t}=0$ and $q^{x}\not=0$}\,,\\ 0\,,\quad&\textrm{ if $(q^{x},q^{t})=(0,0)$}\,.\end{cases}$ (5.9)
@@
where $|Dv|$ is the total variation of $Dv$ and $\frac{Dv}{|Dv|}$ denotes the Radon–Nikodým derivative of $Dv$ with respect to $|Dv|$. For the definition of convex functions of measures, see e.g. [AmFuPa, Section 2.6]. We then consider the auxiliary energy
@@ 5.3 Construction of the inner approximating sequence in a reduced setting
$-\alpha^{\prime}(t)\geq\int_{u^{-1}(t)}\frac{f(z,u(z),\nabla u(z))+|\nabla u(z)|^{p}}{|\nabla u(z)|}{\,{\rm d}}{\mathscr{H}}^{N-1}(z)\,.$ (5.15)
@@
By standard properties of convolution, the family $(v_{\varepsilon,\delta})_{\varepsilon}$ converges to $v_{0,\delta}$ in $L^{1}_{\rm loc}(\Omega\times{\mathbb{R}})$ when ${\varepsilon}\to 0$. Moreover, $Dv_{{\varepsilon},\delta}=Dv_{0,\delta}*_{x}{\varrho}_{\varepsilon}$ for every ${\varepsilon}\in(0,{\varepsilon}_{0}),\delta\in(0,\infty)$, and the family $(Dv_{{\varepsilon},\delta})_{\varepsilon}$ converges weakly-* to $Dv_{0,\delta}$ when ${\varepsilon}\to 0$. From [AmFuPa, Corollary 1.60], this implies that
@@
For every Borel set $A\subset\Omega$ such that ${|A|}=0$, one has $|Dv_{{\varepsilon},\delta}|(A\times{\mathbb{R}})=0$ for ${\varepsilon}\in[0,{\varepsilon}_{0}),\,\delta\in{[0,\infty)}$.
@@
Since $Dv_{0,\delta}$ is the sum of the two mutually singular measures $Dv$ and $(0,\delta\alpha^{\prime}){{\mathscr{H}}^{N+1}}$, one has by [AmFuPa, Proposition 2.37],
@@ 6 The liminf estimate
Suppose $f$ satisfies ($f^{\rm red}$). Let $u\in{\mathscr{E}}(\Lambda)\cap L^{\infty}(\Lambda)$ and let us recall $v_{{\varepsilon},\delta}$ and $q_{{\varepsilon},\delta}$ given by (5.17) and (5.18). Let $\theta:{\mathbb{R}}\to[0,1]$ be a continuous function which vanishes on $[c,\infty)$ for some $c<0$. Then, for every $\delta\in(0,\infty)$, it holds
@@
Let $\eta:{\mathbb{R}}\to[0,1]$ be a continuous compactly supported function such that $\eta\equiv 1$ on $[-M,M]$. Let us define the function $h:\Omega\times{\mathbb{R}}\times\mathbb{S}^{N}\to{\mathbb{R}}$ via
@@
Since $Dv_{{\varepsilon},\delta}$ weakly-* converges to $Dv_{0,\delta}$ when ${\varepsilon}\to 0$ and using also (5.25), we can rely on the Reschetnyak continuity theorem (see e.g. [AmFuPa, Theorem 2.39]) to obtain
@@
$\limsup_{{\varepsilon}\to 0}\left|\int_{\Omega\times{\mathbb{R}}}h\left(x,t,\frac{Dv_{{\varepsilon},\delta}}{|Dv_{{\varepsilon},\delta}|}\right){\,{\rm d}}|Dv_{{\varepsilon},\delta}|-\right.$ $\left.\int_{\Omega\times{\mathbb{R}}}h\left(x,t,\frac{Dv_{0,\delta}}{|Dv_{0,\delta}|}\right){\,{\rm d}}|Dv_{0,\delta}|\right|\leq 4\|h\|_{L^{\infty}}|Dv_{0,\delta}|((\Omega\setminus K_{\tau})\times{\mathbb{R}})$
@@
where the last inequality follows from (6.1) applied to $A=\Omega\setminus K_{\tau}$. In the left-hand side, we rely on the definition of $h$, the fact that $\eta\equiv 1$ on $[-M,M]$ as well as Lemma 5.6 to obtain:
@@ 7.1 Consequences for $\widehat{f}$ of ($H^{\textrm{conv}}$) imposed on $f$
which implies the desired result when $q^{t}<0$. For $q^{t}>0$ or $q^{t}=0$ and $q^{x}\not=0$, the right-hand side of (7.4) is +$\infty$+ ∞. Finally, when $(q^{x},q^{t})=(0,0)$, both sides of (7.4) vanish. ∎
@@
For a.e. $y\in B(x,{\varepsilon})$ and for every $(t,q)\in{\mathbb{R}}\times\mathbb{S}^{N}$, it holds that $F^{{\varepsilon}}(x,t,q)\leq\widehat{f}(y,t,q)$. This inequality remains true for $|{\mathsf{v}}_{t}|\otimes{\mathsf{w}}$ a.e. $(y,t)\in B(x,{\varepsilon})\times{\mathbb{R}}$ and every $q\in\mathbb{S}^{N}$, in view of the fact that $|{\mathsf{v}}_{t}|\otimes{\mathsf{w}}(A\times{\mathbb{R}})=|Dv|(A\times{\mathbb{R}})=0$ for every Borel set $A\subset\Lambda$ such that ${|A|}=0$, see Remark 5.5. It follows that for ${\mathsf{w}}$ a.e. $t\in{\mathbb{R}}$ and every $x\in\Omega$,
@@
We thus recover the Jensen inequality for measures formulated in [Bousquet-Pisa, Proposition 3.11].
@@
where $\eta$ is any function in $C^{\infty}_{c}({\mathbb{R}})$ such that $\eta\equiv 1$ on $[-M,M]$ (recall that $M=\|u\|_{L^{\infty}}$), one gets
@@ 7.2 A key estimate
Using that $u^{-1}(t)$ is ${\mathscr{H}}^{N-1}$ measurable, one has (see [Evans-Gariepy, Section 2.3]),
@@
Indeed, since $q^{t}_{{\varepsilon},\delta}<0$ a.e. for every ${\varepsilon},\delta\geq 0$, one has $\widehat{1}(Dv_{{\varepsilon},\delta})=|\partial_{t}v_{{\varepsilon},\delta}|$. Moreover, we know by (5.25), that $|Dv_{{\varepsilon},\delta}|(\Omega\times{\mathbb{R}})$ converges to $|Dv_{0,\delta}|(\Omega\times{\mathbb{R}})$. Hence, (7.26) follows from the Reshetnyak Continuity Theorem [AmFuPa, Theorem 2.39] applied to the continuous bounded map $q\in\mathbb{S}^{N}\mapsto\beta(q^{t})|q^{t}|$.
@@
where the right-hand side has to be understood as 00 when $\zeta_{\varepsilon}(x,t)=0$. Since $\beta$ is non-decreasing, this gives for $|Dv_{{\varepsilon},\delta}|$ a.e. $(x,t)\in\Omega\times I$,
@@ 8.1 The inner approximation
The ultimate goal of this section is to derive from Corollary 8.1 a suitable approximation of $u$, see Proposition 8.6 below. To this aim, we rely on some technical lemmas from [Bousquet-Pisa]. Let $\Lambda^{\prime}$ be an open subset of $\Lambda$ and $w:\Lambda^{\prime}\times{\mathbb{R}}\to{\mathbb{R}}$ be a bounded Borel map such that $Dw$ is a finite ${\mathbb{R}}^{N+1}$-valued measure and for every $x\in\Lambda^{\prime}$, $t\mapsto w(x,t)$ is non-increasing and left-continuous. We further assume that for every $s\in(0,1)$, there exists $M_{s}>0$ such that for all $t\geq M_{s}$ it holds
@@
Recall the definition of the generalized inverse with respect to the second variable $w^{-1}(\cdot,s)$ introduced in (3.2). For a.e. $s\in(0,1)$, the function $w^{-1}(\cdot,s)$ belongs to $BV(\Lambda^{\prime})\cap L^{\infty}(\Lambda^{\prime})$ and satisfies $\|w^{-1}(\cdot,s)\|_{L^{\infty}(\Lambda^{\prime})}\leq M_{s}$ while the indicator function $\chi_{[w>s]}$ of the set $[w>s]=\{(x,t)\in\Lambda^{\prime}\times{\mathbb{R}}:w(x,t)>s\}$ agrees with $1_{w^{-1}(\cdot,s)}$ a.e. on $\Lambda^{\prime}\times{\mathbb{R}}$, see [Bousquet-Pisa, Lemmas 2.4 and 2.5]. Let $f:\Lambda^{\prime}\times{\mathbb{R}}\times{\mathbb{R}}^{N}\to{[0,\infty)}$ satisfy ($f^{\rm red}$) on $\Lambda^{\prime}$. If $\widehat{E}_{\Lambda^{\prime}\times{\mathbb{R}}}(w)<\infty$, then for a.e. $s\in(0,1)$, the map $w^{-1}(\cdot,s)$ belongs to $W^{1,1}(\Lambda^{\prime})$ while $s\mapsto\widehat{E}_{\Lambda^{\prime}\times{\mathbb{R}}}(\chi_{[w>s]})$ is measurable, with
@@
The last assertion follows from [Bousquet-Pisa, Lemmas 2.6 and 8.5]. In spite of the fact that in the latter reference, those results are stated for a Lagrangian which does not depend on $x$, the extension to the $x$ dependent case is straightforward.
The following abstract fact on the convergence follows from [Bousquet-Pisa, Lemmas 2.9 and 2.10].
@@
For every $0<{\varepsilon}<{\varepsilon}_{0}$ and every $10<\delta<$ 1, the map $v_{{\varepsilon},\delta}$ from (5.17) satisfies (8.1). Moreover,
@@
Observe that $\alpha$ is bounded as any absolutely continuous function, and remember that $\lim_{t\to+\infty}\alpha(t)=0$. Fix $s\in(0,1)$. There exists $M_{s}\geq M=\|u\|_{L^{\infty}(\Lambda)}$ such that for every $t\geq M_{s}$, one has $\alpha(t)<s$. Since $11_{u}(x,t)=$ 1 for every $t\leq-M$ and a.e. $x\in\Lambda$, it follows that $v_{{\varepsilon},\delta}(x,t)=1+\delta\alpha(t)>1$ for all $(x,t)\in\Lambda^{\prime}\times(-\infty,-M]$. Using next that $01_{u}(x,t)=$ 0 for every $t\geq M$, we get $v_{{\varepsilon},\delta}(x,t)=\delta\alpha(t)\leq\alpha(t)<s$ for all $(x,t)\in\Lambda^{\prime}\times[M_{s},\infty)$. This proves that $v_{{\varepsilon},\delta}$ satisfies (8.1) with this value of $M_{s}$, from which (8.5) follows. ∎
To prove that $u_{{\varepsilon},\delta}^{s}$ is Lipschitz, we need the following fact which readily follows from [Bousquet-Pisa, Lemma 2.8].
@@
Since $v_{{\varepsilon},0}(y,t)\geq 0$ and $\alpha$ is decreasing, this implies (8.7) in this case. Similarly, if $t<-M$, then
@@
$-M\leq t\leq t+\tfrac{\overline{C}_{\delta}}{{\varepsilon}}|x-y|\leq M\,.$ (8.9)
@@
Given any $(\delta_{n})_{n\in{\mathbb{N}}}$ decreasing to 00, one can find a sequence $({\varepsilon}_{n})_{n\in{\mathbb{N}}}$ also decreasing to 00 and such that

### v3

@@ 4 Between balance conditions ($H^{\textrm{iso}}$) and ($H^{\textrm{conv}}$)
In particular, $^{w}(s)+s^{\max(p,N)}\leq{\mathcal{C}}_{\widehat{\mathcal{k}}}\mathcal{k}_{2}{\varepsilon}^{-N}+{\mathcal{C}}_{\widehat{\mathcal{k}}}$. Without loss of generality, we may assume that ${\varepsilon}\leq\text{diam}(\Omega)$ and that $\mathcal{k}_{2}$ is sufficiently large to ensure that $\mathcal{k}_{2}\text{diam}(\Omega)^{-N}\geq 1$. Therefore, we have ${w}(s)+s^{\max(p,N)}\leq 2{\mathcal{C}}_{\widehat{\mathcal{k}}}\mathcal{k}_{2}{\varepsilon}^{-N}$. Denoting $\widetilde{\mathcal{k}}=(\mathcal{k}_{1},2{\mathcal{C}}_{\widehat{\mathcal{k}}}\mathcal{k}_{2})$, we get from ($H^{\textrm{iso}}$) that ${w}_{x}(s)\leq{\mathcal{C}}_{\widetilde{\mathcal{k}}}\left({w}(s)+1\right)$, which by (4.4) yields
@@
[…]
@@ 5.2 The convex extension $\widehat{E}$ of the energy $E$
In order to define $\widehat{E}_{\Omega\times{\mathbb{R}}}$, we need to introduce the $^\widehat{\ }$-operation, which goes back to [27] in the context of integral representation of $\Gamma$-limits of variational integrals. For any non-negative convex function $h:{\mathbb{R}}^{N}\to{[0,\infty)}$, we define its recession function $h^{\infty}:{\mathbb{R}^{N}}\to[0,\infty]$ as
@@
It follows from the convexity of $h$ that $\widehat{h}$ is convex on ${\mathbb{R}}^{N+1}$ and positively homogeneous of degree $1$, see e.g. [16, Lemma 8.1]. In particular, for $h\equiv{1}$, we get for every $q=(q^{x},q^{t})\in{\mathbb{R}}^{N}\times{\mathbb{R}}$ with $q^{t}\leq 0$,
@@
[…]
@@
[…]
@@
where $|Dv|$ is the total variation of $Dv$ and $\frac{Dv}{|Dv|}$ denotes the Radon–Nikodým derivative of $Dv$ with respect to $|Dv|$. For the definition of convex functions of measures, see e.g. [2, Section 2.6]. We then consider the auxiliary energy
@@ 5.3 Construction of the inner approximating sequence in a reduced setting
$\alpha^{\prime}(t)\geq\int_{u^{-1}(t)}\frac{f(z,u(z),\nabla u(z))+|\nabla u(z)|^{p}}{|\nabla u(z)|}{\,{\rm d}}{\mathscr{H}}^{N-1}(z)\,.$ (5.15)
@@
By standard properties of convolution, the family $(v_{\varepsilon,\delta})_{\varepsilon}$ converges to $v_{0,\delta}$ in $L^{1}_{\rm loc}(\Omega\times{\mathbb{R}})$ when ${\varepsilon}\to 0$. Moreover, $Dv_{{\varepsilon},\delta}=Dv_{0,\delta}*_{x}{\varrho}_{\varepsilon}$ for every ${\varepsilon}\in(0,{\varepsilon}_{0}),\delta\in(0,\infty)$, and the family $(Dv_{{\varepsilon},\delta})_{\varepsilon}$ converges weakly-* to $Dv_{0,\delta}$ when ${\varepsilon}\to 0$. From [2, Corollary 1.60], this implies that
@@
For every Borel set $A\subset\Omega$ such that |A|=0{|A|}=0, one has $|Dv_{{\varepsilon},\delta}|(A\times{\mathbb{R}})=0$ for ${\varepsilon}\in[0,{\varepsilon}_{0}),\,\delta\in{[0,\infty)}$.
@@
Since $Dv_{0,\delta}$ is the sum of the two mutually singular measures $Dv$ and $(0,\delta\alpha^{\prime}){{\mathscr{H}}^{N+1}}$, one has by [2, Proposition 2.37],
@@ 6 The liminf estimate
Suppose $f$ satisfies ($f^{\rm red}$). Let $u\in{\mathscr{E}}(\Lambda)\cap L^{\infty}(\Lambda)$ and let us recall $v_{{\varepsilon},\delta}$ and $q_{{\varepsilon},\delta}$ given by (5.17) and (5.18). Let $\theta:{\mathbb{R}}\to[0,1]$ be a continuous function which vanishes on $c,\infty)$ for some $c<0$. Then, for every $\delta\in(0,\infty)$, it holds
@@
Let $\eta:{\mathbb{R}}\to[0,1]$ be a continuous compactly supported function such that $\eta\equiv 1$ on [M,M][-M,M]. Let us define the function $h:\Omega\times{\mathbb{R}}\times\mathbb{S}^{N}\to{\mathbb{R}}$ via
@@
Since $Dv_{{\varepsilon},\delta}$ weakly-* converges to $Dv_{0,\delta}$ when ${\varepsilon}\to 0$ and using also (5.25), we can rely on the Reschetnyak continuity theorem (see e.g. [2, Theorem 2.39]) to obtain
@@
[…]
@@
where the last inequality follows from (6.1) applied to $A=\Omega\setminus K_{\tau}$. In the left-hand side, we rely on the definition of $h$, the fact that $\eta\equiv 1$ on [M,M][-M,M] as well as Lemma 5.6 to obtain:
@@ 7.1 Consequences for $\widehat{f}$ of ($H^{\textrm{conv}}$) imposed on $f$
which implies the desired result when $q^{t}<0$. For $q^{t}>0$ or $q^{t}=0$ and $q^{x}\not=0$, the right-hand side of (7.4) is $\infty$. Finally, when $(q^{x},q^{t})=(0,0)$, both sides of (7.4) vanish. ∎
@@
For a.e. $y\in B(x,{\varepsilon})$ and for every $(t,q)\in{\mathbb{R}}\times\mathbb{S}^{N}$, it holds that $F^{{\varepsilon}}(x,t,q)\leq\widehat{f}(y,t,q)$. This inequality remains true for $|{\mathsf{v}}_{t}|\otimes{\mathsf{w}}$ a.e. $(y,t)\in B(x,{\varepsilon})\times{\mathbb{R}}$ and every $q\in\mathbb{S}^{N}$, in view of the fact that $|{\mathsf{v}}_{t}|\otimes{\mathsf{w}}(A\times{\mathbb{R}})=|Dv|(A\times{\mathbb{R}})=0$ for every Borel set $A\subset\Lambda$ such that |A|=0{|A|}=0, see Remark 5.5. It follows that for $\mathsf{w}$} a.e. $t\in{\mathbb{R}}$ and every $x\in\Omega$,
@@
We thus recover the Jensen inequality for measures formulated in [16, Proposition 3.11].
@@
where $\eta$ is any function in $C^{\infty}_{c}({\mathbb{R}})$ such that $\eta\equiv 1$ on [M,M][-M,M] (recall that $M=\|u\|_{L^{\infty}}$), one gets
@@ 7.2 A key estimate
Using that $u^{-1}(t)$ is ${\mathscr{H}}^{N-1}$ measurable, one has (see [33, Section 2.3]),
@@
Indeed, since $q^{t}_{{\varepsilon},\delta}<0$ a.e. for every ${\varepsilon},\delta\geq 0$, one has $\widehat{1}(Dv_{{\varepsilon},\delta})=|\partial_{t}v_{{\varepsilon},\delta}|$. Moreover, we know by (5.25), that $|Dv_{{\varepsilon},\delta}|(\Omega\times{\mathbb{R}})$ converges to $|Dv_{0,\delta}|(\Omega\times{\mathbb{R}})$. Hence, (7.26) follows from the Reshetnyak Continuity Theorem [2, Theorem 2.39] applied to the continuous bounded map $q\in\mathbb{S}^{N}\mapsto\beta(q^{t})|q^{t}|$.
@@
where the right-hand side has to be understood as 0 when $\zeta_{\varepsilon}(x,t)=0$. Since $\beta$ is non-decreasing, this gives for $|Dv_{{\varepsilon},\delta}|$ a.e. $(x,t)\in\Omega\times I$,
@@ 8.1 The inner approximation
The ultimate goal of this section is to derive from Corollary 8.1 a suitable approximation of $u$, see Proposition 8.6 below. To this aim, we rely on some technical lemmas from [16]. Let $\Lambda^{\prime}$ be an open subset of $\Lambda$ and $w:\Lambda^{\prime}\times{\mathbb{R}}\to{\mathbb{R}}$ be a bounded Borel map such that $Dw$ is a finite ${\mathbb{R}}^{N+1}$-valued measure and for every $x\in\Lambda^{\prime}$, $t\mapsto w(x,t)$ is non-increasing and left-continuous. We further assume that for every $s\in(0,1)$, there exists $M_{s}>0$ such that for all $t\geq M_{s}$ it holds
@@
Recall the definition of the generalized inverse with respect to the second variable $w^{-1}(\cdot,s)$ introduced in (3.2). For a.e. $s\in(0,1)$, the function $w^{-1}(\cdot,s)$ belongs to $BV(\Lambda^{\prime})\cap L^{\infty}(\Lambda^{\prime})$ and satisfies $\|w^{-1}(\cdot,s)\|_{L^{\infty}(\Lambda^{\prime})}\leq M_{s}$ while the indicator function $\chi_{[w>s]}$ of the set $w>s]=\{(x,t)\in\Lambda^{\prime}\times{\mathbb{R}}:w(x,t)>s\}$ agrees with $1_{w^{-1}(\cdot,s)}$ a.e. on $\Lambda^{\prime}\times{\mathbb{R}}$, see [16, Lemmas 2.4 and 2.5]. Let $f:\Lambda^{\prime}\times{\mathbb{R}}\times{\mathbb{R}}^{N}\to{[0,\infty)}$ satisfy ($f^{\rm red}$) on $\Lambda^{\prime}$. If $\widehat{E}_{\Lambda^{\prime}\times{\mathbb{R}}}(w)<\infty$, then for a.e. $s\in(0,1)$, the map $w^{-1}(\cdot,s)$ belongs to $W^{1,1}(\Lambda^{\prime})$ while $s\mapsto\widehat{E}_{\Lambda^{\prime}\times{\mathbb{R}}}(\chi_{[w>s]})$ is measurable, with
@@
The last assertion follows from [16, Lemmas 2.6 and 8.5]. In spite of the fact that in the latter reference, those results are stated for a Lagrangian which does not depend on $x$, the extension to the $x$ dependent case is straightforward.
The following abstract fact on the convergence follows from [16, Lemmas 2.9 and 2.10].
@@
For every $0<{\varepsilon}<{\varepsilon}_{0}$ and every $0<\delta<1$, the map $v_{{\varepsilon},\delta}$ from (5.17) satisfies (8.1). Moreover,
@@
Observe that $\alpha$ is bounded as any absolutely continuous function, and remember that $\lim_{t\to+\infty}\alpha(t)=0$. Fix $s\in(0,1)$. There exists $M_{s}\geq M=\|u\|_{L^{\infty}(\Lambda)}$ such that for every $t\geq M_{s}$, one has $\alpha(t)<s$. Since $1_{u}(x,t)=1$ for every $t\leq-M$ and a.e. $x\in\Lambda$, it follows that $v_{{\varepsilon},\delta}(x,t)=1+\delta\alpha(t)>1$ for all $(x,t)\in\Lambda^{\prime}\times(-\infty,-M]$. Using next that $1_{u}(x,t)=0$ for every $t\geq M$, we get $v_{{\varepsilon},\delta}(x,t)=\delta\alpha(t)\leq\alpha(t)<s$ for all $(x,t)\in\Lambda^{\prime}\times[M_{s},\infty)$. This proves that $v_{{\varepsilon},\delta}$ satisfies (8.1) with this value of $M_{s}$, from which (8.5) follows. ∎
To prove that $u_{{\varepsilon},\delta}^{s}$ is Lipschitz, we need the following fact which readily follows from [16, Lemma 2.8].
@@
Since $v_{{\varepsilon},0}(y,t)\geq 0$ and $\alpha$ is decreasing, this implies (8.7) in this case. Similarly, if t<Mt<-M, then
@@
$M\leq t\leq t+\tfrac{\overline{C}_{\delta}}{{\varepsilon}}|x-y|\leq M\,.$ (8.9)
@@
Given any $(\delta_{n})_{n\in{\mathbb{N}}}$ decreasing to 0, one can find a sequence $({\varepsilon}_{n})_{n\in{\mathbb{N}}}$ also decreasing to 0 and such that